OpenStudy (anonymous):

The ration of country albums to jazz albums in a music collection is 2:3. If the music collection has 45 albums, how many are country albums? Would it be 2x*3x=45? Then would you do 2x*3x and get 6x=45. Then divide by 6 and get 7.5? I don't think this is right, help!

OpenStudy (raden):

2/(2+3) * 45 = 2/5 * 45 = ...

OpenStudy (anonymous):

so I write that? and solve 2/5*45?

OpenStudy (raden):

yup for be easier u can do 45/5 first then times by 2

OpenStudy (raden):

got it ?

OpenStudy (anonymous):

no

OpenStudy (raden):

dont be shy, use the calculator :P

OpenStudy (anonymous):

so i type 45/5 on the calcultor and get 225, then divide that by 2?

OpenStudy (raden):

ur calculator was low-battery or be error :P 45/5 = 9 not 225

OpenStudy (anonymous):

I typed it wrong oops

OpenStudy (anonymous):

my bad

OpenStudy (raden):

hihi :)

OpenStudy (anonymous):

then you get 9/2 which is 4.5?

OpenStudy (raden):

9 times 2 not divided

OpenStudy (anonymous):

so you get 18?

OpenStudy (raden):

yup

OpenStudy (anonymous):

how do I write this on my paper?

OpenStudy (raden):

2/(2+3) * 45 = 2/5 * 45 = 18

OpenStudy (anonymous):

Thanks!(:

OpenStudy (raden):

very welcome :)